# Integral Representations of Three Novel Multiple Zeta Functions for Barnes Type: A Probabilistic Approach

### 13 Dec 2023 v1

@@
positive real numbers. Denote
and As usual, we
adopt the notations and for real
@@
It is seen that the function in (3) above
generalizes the Barnes multiple zeta function (see
Corollary 1 below). We will provide the integral representations of
the above three novel multiple zeta functions through the hyperbolic
probability distributions given by Pitman and Yor (2003), and
investigate their analytically continued properties including the
set of singularities. Surprisingly, unlike the previous results, the
functions and in (4) and
(5) above can extend analytically to entire functions in
The main results are stated in Section 3. Section 2 provides the
needed lemmas. Finally, the proofs of the main results are given in
Section 4.
@@ 2. Lemmas
integral theorem (see, e.g., Sharma 2005, Chapter 2). The crucial
Lemma 3 can be proved by Moreraβs theorem (see Cheng and Hu 2022).
The hyperbolic random variables in Lemma 4 are introduced and
@@
be an integer and let be a continuous function on
Then we have
where the parameters and the integral
region
@@
For any positive real define the complex-valued function
Then is an entire function in
Remark 1. Note that for
the function value of
becomes a linear combination of moments of Thus, for
simplicity, we might call the integral transformation a
moment function of the distribution
@@
that , and
are positive real numbers satisfying Denote
@@
multiple hyperbolic sinh-moment function in
has an integral representation of the form:
@@
The zeta function in can extend
@@
function in is an analytic function in
except for Precisely,
@@
function in has trivial zeros at
and
@@
If
then and the multiple hyperbolic sinh-moment function
in (3) reduces to the Barnes multiple zeta
function Thus, we have the following.
@@
where
and are independent
random variables, each having the
density function
@@
hyperbolic sinh-moment function reduces to the
Hurwitz zeta function Thus, we have the following.
@@
and the multiple hyperbolic sinh-moment function
in (3) reduces to
@@
where has the density
function
@@
multiple hyperbolic cosh-moment function in
has an integral representation of the form:
@@
function in can extend analytically to
@@
and the multiple hyperbolic cosh-moment function
in (4) reduces to the following zeta function
@@
where has the density function in Moreover,
the zeta function in extends to an entire
function in
@@
and the multiple hyperbolic cosh-moment function
in (4) reduces to
@@
where has the density
function
@@
tanh-moment function in has an integral
representation of the form:
@@
function in can extend analytically to an
entire function in
@@
the multiple hyperbolic tanh-moment function in
(5) reduces to
@@
In these cases, Lemma 2 applies. The results
are too complicated to be presented here.
@@ 4. Proofs of the main results
multiple hyperbolic sinh-moment function in (3)
by we have
@@
and the binomial series
or, equivalently,
@@
due to Lemma 3. Therefore, the zeta function in
(6) can extend analytically to a meromorphic function in
and all the poles are simple due to Lemma 1. This proves part (a).
Parts (b) and (c) follow from part (a) and Lemma 1 again
@@
relationship between and
specifically, we can rewrite
@@
sinh-moment function in (4) by we
have
@@
Lemma 3. Therefore, the zeta function in (12)
can extend analytically to an entire function in This
completes the proof of Theorem 2.
@@
Multiplying both sides of the function in (5)
by we have

### 02 Jan 2024 v2

@@
positive real numbers. Denote and
As usual, we adopt the
notations and for real
@@
It is seen that when all the function
in (3) above reduces to the Barnes
multiple zeta function (see Corollary 1 below). We will
provide the integral representations of the above three novel
multiple zeta functions through the hyperbolic probability
distributions given by Pitman and Yor (2003), and investigate their
analytically continued properties including the set of
singularities. Surprisingly, unlike the previous results, the
functions and in (4) and (5) above can extend analytically to
entire functions in The main results are stated in Section
3. Section 2 provides the needed lemmas. Finally, the proofs of the
main results are given in Section 4.
@@ 2. Lemmas
integral theorem (see, e.g., Sharma 2005, Chapter 2, or Gibson
1931, p.β492). The crucial Lemma 3 modifies slightly
Cheng and Huβs (2022) Lemma 2.1. The
hyperbolic random variables in Lemma 4 are introduced and
@@
be an integer and let be a Lebesgue measurable function
on Then we have
provided that the resulting single integral on the right exists.
Here, the parameters and the region of
multiple integral
@@
For any nonzero real define the complex-valued function
where Then is an
entire function in
Proof. We sketch the proof below. For each define the
closed square
Denote the integer Then there exists a constant
such that
This implies that the integral transformation is
absolutely convergent and converges uniformly in each compact
subset of Moreover, is a continuous function in
Let be any closed triangle in and recall
that for each fixed real is an entire function in
Then we have, by Cauchyβs theorem,
where is the boundary of Finally,
Moreraβs theorem completes the proof.
Remark 1. Note that for the
function value of becomes a linear combination of moments
of Thus, for simplicity, we might call the integral
transformation a moment function of the distribution
@@
that , and are
positive real numbers satisfying Denote
@@
multiple hyperbolic sinh-moment function in has an integral representation of the form:
@@
The zeta function in
can extend
@@
function in is an analytic
function in except for Precisely,
@@
function in has trivial zeros
at and
@@
If then
and the multiple hyperbolic sinh-moment function in (3) reduces to the Barnes multiple zeta function
Thus, we have the following.
@@
where and are
independent random variables, each
having the density function
@@
hyperbolic sinh-moment function
reduces to the Hurwitz zeta function Thus, we have the
following.
@@
and the multiple hyperbolic sinh-moment function in (3) reduces to
@@
where and has
the density function
@@
multiple hyperbolic cosh-moment function in has an integral representation of the form:
@@
function in can extend
analytically to
@@
and the multiple hyperbolic cosh-moment function in (4) reduces to the following zeta function
@@
where and has the density function
in Moreover, the zeta function in
extends to an entire function in
@@
and the multiple hyperbolic cosh-moment function in (4) reduces to
@@
where and has
the density function
@@
tanh-moment function in has
an integral representation of the form:
@@
function in can extend
analytically to an entire function in
@@
the multiple hyperbolic tanh-moment function in (5) reduces to
@@
In these cases,
Lemma 2 applies. The results are too complicated to be presented
here.
@@ 4. Proofs of the main results
multiple hyperbolic sinh-moment function in (3) by we have
@@
and the binomial series or, equivalently,
@@
due to Lemma 3. Therefore, the zeta function in (6) can extend analytically to a meromorphic function
in and all the poles are simple due to Lemma 1. This proves
part (a). Parts (b) and (c) follow from part (a) and Lemma 1 again
@@
relationship between and specifically, we can rewrite
@@
sinh-moment function in (4) by
we have
@@
Lemma 3. Therefore, the zeta function in (12) can extend analytically to an entire function in
This completes the proof of Theorem 2.
@@
Multiplying both sides of the function in (5) by we have